A bomber flies horizontally with a speed of 347 m/s relative to the ground. The altitude of the bomber is 2900 m and the terrain is level. Neglect the effects of air resistance.
The acceleration of gravity is 9.8 m/s2 .
a) How far from the point vertically un- der the point of release does a bomb hit the ground?
Answer in units of m.

b) At what angle from the vertical at the point of release must the telescopic bomb sight be set so that the bomb hits the target seen in the sight at the time of release?
Answer in units of ◦.
Answer #1

a) time take by bomb to reach ground,

vertically;

2900 = 0 + 9.8t2 /2
t = 24.32 sec
d = ut = 347 * 24.32 = 8439.04 m

b) angle from vertical = tan-1(8439.04 / 2900) = 71.03 degrees